4) Describe how you would prepare 1.0 L of a 0.10 M solution of sulfuric acid from a 3.0 M solution of sulfuric acid.
5) Describe how you would prepare 500 mL of a 1.0 M solution of potassium chloride that is 93.0% pure. The gram formula weight of potassium chloride is 74.56 g/mol.
6) Describe how you would prepare 1000 mL of a 1.0 N solution of magnesium hydroxide. The gram formula weight of magnesium hydroxide is 58.33 g/mo
1
Expert's answer
2020-07-22T05:48:39-0400

4) C1V1 = C2V2 formula can be used to find the volume of the initial solution:

V1 = C2V2 / C1 = (1.0L x 0.10M) / 0.30M = 0.033 L

Therefore, 0.033 L or 33 mL of 3.0 M solution should be measured using the graduated cylinder, moved to the 1.0 L volumetric flask and diluted with water up to the 1.0 L mark.


5) 500 mL of 1.0 M KCl contains 1.0 x 0.5 = 0.5 moles of pure KCl.

m (KCl) = 0.5 mol x 74.56 g/mol = 37.28 g.

Considering the 93% purity of the sample, the actual mass weighted for preparation equals 37.28g / 0.93 = 40.09 g. Therefore, 40.09 g of the impure KCl should be weighted using the electronic balance, placed into the graduated glassware and water up to 500 mL mark should be added.


6) Assuming that Mg(OH)2 is ideally soluble in water, the normality for this compound equals molarity x 2 because one mole of Mg(OH)2 produces 2 moles of OH- ions. Therefore the molarity of the solution is 1.0 / 2 = 0.5 M.

1000 mL of the solution contain 0.5 moles of Mg(OH)2.

m (Mg(OH)2) = 0.5 mol x 58.33 g/mol = 29.17 g.

So 29.17 g of Mg(OH)2 should be weighted using the electronic balance, placed into the graduated glassware and water up to 1000 mL mark should be added.